OpenStudy (anonymous):

What is the fraction 8/36 expressed in lowest terms?

OpenStudy (anonymous):

giving medals

OpenStudy (amorfide):

think of it as a ration for every 8 of bananas, you have 36 apples 8:36 what number is a factor of 8 and 36 list factors of 8 1,2,4,8 list factors of 36 1,2,3,4,6,12,18 we take the highest factor that is common in both, in this case 4 so our highest factor of each number is 4 we can take out a factor of 4 from 8 and 36 to get \[\frac{ 8 }{ 36 }= \frac{ 4 \times 2 }{ 4 \times 9 }\] the 4's will become 1 that leaves us with \[\frac{ 2 }{ 9 }\]

OpenStudy (amorfide):

ratio*